OpenStudy (anonymous):

"A soda machine contains 20 coins. Some of the coins are nickels and the rest are quarters. The value of the coins is $4.40. Find the number of coins of each kind

OpenStudy (anonymous):

n+q=20 5n+25q=440 (mult first one by five) 5n+5q=100 (subtract) 20q=340 q=17 so 17 quarters and 3 nickels